Hello, I appeciate very much that there are people like you up there; willin and dedicated to help people like me. Thank you very much for your effort.
My dilemma is and have always been word problems. They are not like other math problems where you just follow or apply a given formula. Words problems are always changin and I never know when to use what. I've always had difficulty translating words to the actual equation which by the way I have no problem solving. If there is some guidelines you can show me as to when I use what, I would appreciate it very, very much. The problems are:
1) Henry has $1.35 in nickels and dimes. There are 9 more nickels than dimes. find the number he has of each kind.
2) Mark has $5.05 in quarters and dimes. The number of quarters exceeds twice the number of dimes by 1. Find the number he has of each.
raulmiller@optonline.net

Answer by ewatrrr(24785)


Hi,
Note***convert these problems to how many pennys or cents (they are all done that way)
1)Let x represent the number of dimes, then (x+9)is the number of nickels
5(x+9) + 10x = 135cents
solve for x
2) Let x represent the number of dimes, then (2x+1)is the number of quarters
10x + 25(2x+1) = 505 cents
solve for x
60x + 25 = 505
60x = 480
x = 8 dimes and therefore (2*8+1) quarters Or 17 quarters
checking our answer
80 + 425 = 505cents Or $5.05